# Language-Based Bayesian Optimization Research Assistant (BORA)

### v1

@@ A.0.3 Ablation Studies
In this section, we assess BORA’s flexibility in dynamically switching between LLM-dominated search strategies and BO-dominated ones in the optimization. A key parameter that regulates this is the plateau window size $[m_{\text{min}}=0,\quad m_{\text{max}}=3m_{\text{init}}]$, conditioned by the chosen initial value of the plateau $m_{\text{init}}$. In this ablation study, we test how different values of $m_{\text{init}}$ affect BORA’s performance.
We select $m_{\text{init}}=2$, which translates into a narrower plateau window of [0, 6]06[0,\ 6][ 0 , 6 ] and thus increases the number of plateaus encountered during the optimization, leading to more frequent LLM interventions. Additionally, we test $m_{\text{init}}=32$, which corresponds to fewer LLM interventions and more BO steps. We test BORA’s performance on both of these scenarios on the Ackley (15D) function. Note that we also include in the comparison:
@@ A.0.4 Cost of BORA Runs
[…]
plus the LLM calls incurred by the reflection strategies. These findings demonstrate that BORA’s strategic use of LLMs, beyond performing significantly better, has much smaller LLM inference and cost footprints.
@@ B.1 Levy Function (10D)
The Levy function Surjanovic and Bingham [2024] is a popular synthetic function used to test global optimization methods, particularly in high-dimensional black-box functions. We restrict its search space to $[-10,10]^{10}$ with a global maximum realized at $f(x)=806.789$ when $x=(1,1,\ldots,1)$.
@@ B.4 Pétanque (7D)
Rainfall mm Total daily precipitation. [0, 20]
@@
[…]
@@ B.6 Hydrogen Production (10D)
The Hydrogen Production experiment was orignally realized by Burger et al. to find the optimal composition of a ten-material mixture maximizing the Hydrogen Evolution Rate (HER) via photocatalysis. Due to physical constraints in the experiment,  Burger et al. constrained the search space to keep the total mixture volume less than 5 mL (the maximum volume in the glass vial). Other constraints come from the liquid dispensing module (minimum accurate dispense = 0.25 mL) and the solid dispensing module (minimum dispense = 0.2 mg), turning the idealized continuous space into a discretized one. The resulting combinatorial search space was extremely large, with 98,423,325 possible combinations, Burger et al., so  Burger et al. used an autonomous mobile robotic chemist along with BO to search for the optimal composition of the chemical mixture.
To replicate this experiment, we used the same cost-effective and time-saving method as Cissé et al. who are part of the same research group as  Burger et al.Cissé et al. built an oracle method by training a Gaussian process regression on a dataset of 1119 experimental observations of the original study to serve as ‘ground truth’. Although the interpolated model is an approximation,  Cissé et al. demonstrated, in their work about HypBO, that it captured the underlying chemical processes of the original experiment, making it reliable enough to draw safe conclusions.
@@
  • Objective: Maximize the HER ($\mu$mol/h), by optimizing the quantities of each of the ten chemicals in the mixture.
@@
  • Constraint: To ensure compatibility with the experimental setup,
@@
Appendix C BORA
BORA’s Vanilla BO action $a_{1}$ is built on top of BayesOpt Nogueira [2014] and uses Expected Improvement (EI) as an acquisition function. The Gaussian Process (GP) employs a Matérn kernel, with hyperparameters automatically optimized based on the experimental data to best fit the model. The default LLM model is OpenAI’s GPT-4o-mini which as empirically selected based on efficiency, performance and cost criteria amongst other available LLMs. However, given that BORA was designed to work with the OpenAI API, it allows for any model from the API.
Appendix D Baselines
Figure 12: LAEA vs. LAEA_5. LAEA denotes LAEA with an initial population size of 15. Solid lines show mean values while shaded areas indicate ±$\pm 0.25$ standard error.
All the baselines we tested BORA against have internal random sampling mechanisms to generate random points. In the BO-based baselines, these random points are partially used as candidate points for the acquisition function. In the LAEA method, they contribute to the generation and evolution of the population during the optimization. To enable a fair comparison between BORA and baseline methods for the Hydrogen Production experiment, we adapted the random sampling mechanism of each baseline to effectively handle the discrete and constrained nature of the input space. This ensures that the baselines remain competitive and adhere to the unique constraints of the Hydrogen Production problem. Note that HypBO did not need modification as it inherently supports such spaces.
@@ Appendix D Baselines
D.1 LAEA Population Size
@@
D.2 ColaBO Input Priors by an LLM
As stated in the manuscript, to avoid the impracticality of relying on humans to provide inputs for multiple trials across all experiments for ColaBO, we used the LLM GPT-4o-mini to generate the ‘human’ inputs. It is possible, of course, that as for ColaBO, human domain experts will in some cases provide superior inputs than produced by an LLM. ColaBO uses GP priors to express human input as:
@@ D.2 ColaBO Input Priors by an LLM
D.3 HypBO Hypotheses by an LLM
Unlike in BORA, HypBO Cissé et al. [2024] defines a hypothesis as a static region of interest; that is, a ‘box soft constraint’. We automate the generation of these hypotheses in testing HypBO on the benchmarks by using an LLM. Similar to the ColaBO input prior generation, we used LLM GPT-4o-mini and the same task description prompts as used for BORA. Again, the quality of the LLM input may not match that of a human domain expert, but this was done because of the impracticality of drawing on human experts over multiple trial optimization runs. The LLM is tasked to predict the location of the optimum and to give a [0, 1] confidence score. Because HypBO supports the use of multiple static hypotheses, we do not restrict the LLM in the number of hypotheses that it can recommend. Then, for each hypothesis, we use the corresponding score to ‘draw’ a hyper-rectangle around its suggested optimum location. The length of the hyper-rectangle $L$ corresponds to the experiment space scaled using the confidence score as follows:
@@ D.3 HypBO Hypotheses by an LLM
D.3.1 Case of the Hydrogen Production Benchmark
@@
Appendix E Extended Methodology
As BORA is a hybrid method involving LLMs, its effectiveness depends in part on the quality and the structure of the LLM outputs. Additionally, ensuring that the LLM outputs respect a desired format guarantees that BORA is fully automated. This section details the LLM prompt engineering, including the reflection strategies used to build robust prompts for the LLM agent to:
@@ Appendix E Extended Methodology
E.1 LLM Role-Based Prompting
@@
E.1.1 Experiment Card
This card is constructed by the user following the standardized template shown in Figure 13. This template allows the construction of a comprehensive description of the experiment, helping the LLM contextualize the problem at hand. Figure 14 shows what that looks like for the Hydrogen Production experiment.
@@ E.1.1 Experiment Card
E.1.2 Enhanced Role Prompting
@@ E.3 Optimization Initialization
E.4 LLM Interventions
E.4.1 $a_{2}$ LLM Comments and Suggests $n_{\text{LLM}}$ Points
@@
E.4.2 $a_{3}$ LLM Comments and Selects 2 BO Points
The LLM is prompted following the template shown in Figure 25 with the gathered data up to the current iteration, a set of 5 BO-suggested points, and any previous comments. Then, it is tasked to comment on the optimization progress in light of the new data and is constrained to select the two most promising BO candidates that best align with its hypotheses for maximizing the target objective.
@@ E.4.2 $a_{3}$ LLM Comments and Selects 2 BO Points
E.5 Optimization Conclusions and Summary
@@ E.5.2 Optimization Summary
E.6 LLM Reflection and Fallback Strategies
@@
E.6.1 Self-consistency
For the LLM interventions $a_{2}$ and $a_{3}$, we leverage self-consistency by generating $n=3$ outputs from the same prompt. Each of those outputs reflects a potential viewpoint. These outputs are then systematically consolidated to synthesize a comprehensive and unified response. This consolidation process involves several key steps: (a) reflecting and critiquing each output, (b) evaluating consistency, and resolving discrpancies by cross-referencing them against the optimization dataset. This self-consistency procedure helps maintain coherence and accuracy, which minimizes the risk of biases from a single output while fostering a more robust and nuanced proposition.
E.6.2 Feasibility Verification and Fallback Mechanisms
@@
While context window limits were a pronounced limitation when LLMs were first introduced, their capabilities here have increased markedly, resulting in much larger context windows. To this point, although we have designed a summarizing solution in case the LLM’s context window is reached, we have not encountered that scenario during our tests here. This could occur with more complex or longer experiments. Our summarizing mechanism works as follows. We track the number of tokens in the LLM agent’s chat history and whether the context window will be exhausted at the subsequent LLM intervention. In such a scenario, instead of giving all the previously gathered data to the LLM as a part of its intervention package, we supply it with the correlation matrix of the data instead. The correlation matrix summarizes the pairwise relationships between the input variables of the problem at hand, considerably reducing the input size for the LLM. In fact, for a dataset with $d$ input variables and $N>>d$ number of samples, the correlation matrix has only $d\times d$ elements instead of the raw dataset with $d\times N$ elements. Although the correlation matrix summarizes the data, the LLM loses the intricate details in the raw data. However, as a result of a condensed data summary, the LLM processes it much faster.

### v2

@@ A.0.3 Ablation Studies
In this section, we assess BORA’s flexibility in dynamically switching between LLM-dominated search strategies and BO-dominated ones in the optimization. A key parameter that regulates this is the plateau window size $m_{\text{min}}=0,\quad m_{\text{max}}=3m_{\text{init}}]$, conditioned by the chosen initial value of the plateau $m_{\text{init}}$. In this ablation study, we test how different values of $m_{\text{init}}$ affect BORA’s performance.
We select $m_{\text{init}}=2$, which translates into a narrower plateau window of [0, 6][0,\ 6] and thus increases the number of plateaus encountered during the optimization, leading to more frequent LLM interventions. Additionally, we test $m_{\text{init}}=32$, which corresponds to fewer LLM interventions and more BO steps. We test BORA’s performance on both of these scenarios on the Ackley (15D) function. Note that we also include in the comparison:
@@ A.0.4 Cost of BORA Runs
We report the total cost of running BORA and LAEA in the manuscript experiment section, both using GPT-4o-mini as the LLM component, which covers 6 benchmarks across 10 trials with a 105-sample budget. The total cost of running BORA amounts to about 5 US Dollars (USD) at today’s prices, while the cost for LAEA with an initial population size of 15 is around 41 USD, considering the LLM costs only. This cost difference arises from the inherent design of the two methods. As an LLM-BO hybrid, BORA alternates between LLM-based and BO-based sampling via its adaptive policy. Notably, the LLM is more involved in optimization when it proves trustworthy. In contrast, LAEA uses the LLM as a surrogate model to evaluate the fitness of the individuals in the population. Given that at each iteration during the optimization, LAEA uses the LLM to estimate the fitness of each individual in the population, this translates into a much more frequent LLM call, leading to increased computational and inference costs. To put this into perspective for a population size of $\text{pop\_size}=15$ individuals, and a maximum number of objective function evaluations of $n_{\text{evals}}=105$, LAEA requires at least
$n_{\text{LLM\_calls}}=(n_{\text{evals}}-\text{pop\_size})\times\text{pop\_size}=1350,$
plus the LLM calls incurred by the reflection strategies. These findings demonstrate that BORA’s strategic use of LLMs, beyond performing significantly better, has much smaller LLM inference and cost footprints, noting that the use of LLMs also has a significant (and growing) energy cost.
@@ B.1 Levy Function (10D)
The Levy function Surjanovic and Bingham [2024] is a popular synthetic function used to test global optimization methods, particularly in high-dimensional black-box functions. We restrict its search space to $10,10]^{10}$ with a global maximum realized at $f(x)=806.789$ when $x=(1,1,\ldots,1)$.
@@ B.4 Pétanque (7D)
Rainfall $m$ Total daily precipitation. [0, 20]
@@
score $=100\cdot\exp(-\text{distance\_error}\cdot 0.2)$ (11)
@@ B.6 Hydrogen Production (10D)
The Hydrogen Production experiment was orignally realized by Burger et al. to find the optimal composition of a ten-component mixture maximizing the Hydrogen Evolution Rate (HER) via photocatalysis. Due to physical constraints in the experiment,  Burger et al. constrained the search space to keep the total mixture volume less than 5 mL (the maximum volume in the reaction vial). Other constraints come from the liquid dispensing module (minimum accurate dispense = 0.25 mL) and the solid dispensing module (minimum dispense = 0.2 mg), turning the idealized continuous space into a discretized one. The resulting combinatorial search space was extremely large, with 98,423,325 possible combinations, Burger et al., so the authors used an autonomous mobile robotic chemist along with BO to search for the optimal composition of the chemical mixture.
To replicate this experiment, we used the same cost-effective and time-saving method as Cissé et al., who are part of the same research group as  Burger et al.Cissé et al. built an oracle method by training a Gaussian process regression on a dataset of 1119 experimental observations of the original study to serve as ‘ground truth’, or oracle model. Although this interpolated model is an approximation,  Cissé et al. demonstrated in their HypBO work that it captured the underlying chemical processes of the original experiment, making it reliable enough to draw safe conclusions.
@@
[…]
@@
  • Constraints: To ensure compatibility with the experimental setup,
@@
Appendix C Baselines
Figure 12: LAEA vs. LAEA_5. LAEA denotes LAEA with an initial population size of 15. Solid lines show mean values while shaded areas indicate $\pm 0.25$ standard error.
All of the baselines that we tested BORA against have internal random sampling mechanisms to generate random points. In the BO-based baselines, these random points are partially used as candidate points for the acquisition function. In the LAEA method, they contribute to the generation and evolution of the population during the optimization. To enable a fair comparison between BORA and baseline methods for the Hydrogen Production experiment, we adapted the random sampling mechanism of each baseline to effectively handle the discrete and constrained nature of the input space. This ensures that the baselines remained competitive and adhered to the unique constraints of the Hydrogen Production problem. Note that HypBO did not need modification as it inherently supports such spaces.
@@ Appendix C Baselines
C.1 LAEA Population Size
@@
C.2 ColaBO Input Priors by an LLM
As stated in the manuscript, to avoid the impracticality of relying on humans to provide inputs for multiple trials across all experiments for ColaBO, we used the LLM GPT-4o-mini to generate the ‘human’ inputs. It is likely, of course, that human domain experts will in some cases provide superior inputs than produced by an LLM, and we note that limitation here. ColaBO uses GP priors to express human input as:
@@ C.2 ColaBO Input Priors by an LLM
C.3 HypBO Hypotheses by an LLM
Unlike in BORA, HypBO Cissé et al. [2024] defines a hypothesis as a static region of interest; that is, a ‘box soft constraint’. We automated the generation of these hypotheses in testing HypBO on the benchmarks by using an LLM. Similar to the ColaBO input prior generation, we used LLM GPT-4o-mini and the same task description prompts as used for BORA. Again, the quality of the LLM input may not match that of a human domain expert, but this was done because of the impracticality of drawing on human experts over multiple trial optimization runs. The LLM is tasked to predict the location of the optimum and to give a [0, 1] confidence score. Because HypBO supports the use of multiple static hypotheses, we do not restrict the LLM in the number of hypotheses that it can recommend. Then, for each hypothesis, we use the corresponding score to ‘draw’ a hyper-rectangle around its suggested optimum location. The length of the hyper-rectangle $L$ corresponds to the experiment space scaled using the confidence score as follows:
@@ C.3 HypBO Hypotheses by an LLM
C.3.1 Case of the Hydrogen Production Benchmark
@@
Appendix D Extended Methodology
As BORA is a hybrid method involving LLMs, its effectiveness depends, in part, on the quality and the structure of the LLM outputs. Additionally, ensuring that the LLM outputs respect a desired format guarantees that BORA is fully automated. This section details the LLM prompt engineering, including the reflection strategies used to build robust prompts for the LLM agent to:
@@ Appendix D Extended Methodology
D.1 LLM Role-Based Prompting
@@
D.1.1 Experiment Card
This card is constructed by the user following the standardized template shown in Figure 13. This template allows the construction of a comprehensive description of the experiment, helping the LLM contextualize the problem at hand. Figure 14 shows what this looks like for the Hydrogen Production experiment.
@@ D.1.1 Experiment Card
D.1.2 Enhanced Role Prompting
@@ D.3 Optimization Initialization
D.4 LLM Interventions
D.4.1 $a_{2}$ LLM Comments and Suggests $n_{\text{LLM}}$ Points
@@
D.4.2 $a_{3}$ LLM Comments and Selects $n_{\text{LBO}}$ BO Points
The LLM is prompted following the template shown in Figure 25 with the gathered data up to the current iteration, a set of $n_{\text{BO}}=5$ BO-suggested points, and any previous comments. Then, it is tasked to comment on the optimization progress in light of the new data and is constrained to select the $n_{\text{LBO}}=2$ most promising BO candidates that best align with its hypotheses for maximizing the target objective.
@@ D.4.2 $a_{3}$ LLM Comments and Selects $n_{\text{LBO}}$ BO Points
D.5 Optimization Conclusions and Summary
@@ D.5.2 Optimization Summary
D.6 LLM Reflection and Fallback Strategies
@@
D.6.1 Self-consistency
For the LLM interventions $a_{2}$ and $a_{3}$, we leverage self-consistency by generating $n=3$ outputs from the same prompt. Each of those outputs reflects a potential viewpoint. These outputs are then systematically consolidated to synthesize a comprehensive and unified response. This consolidation process involves several key steps: (a) reflecting and critiquing each output, (b) evaluating consistency, and resolving discrepancies by cross-referencing them against the optimization dataset. This self-consistency procedure helps maintain coherence and accuracy, which minimizes the risk of biases from a single output while fostering a more robust and nuanced proposition.
D.6.2 Feasibility Verification and Fallback Mechanisms
@@
While context window limits were a pronounced limitation when LLMs were first introduced, their capabilities here have increased markedly, resulting in much larger context windows. This is likely to evolve further. To this point, although we have designed a summarizing solution in case the LLM’s context window is reached, we have not encountered that scenario during our tests here. This could occur with more complex or longer experiments. Our summarizing mechanism works as follows. We track the number of tokens in the LLM agent’s chat history and whether the context window will be exhausted at the subsequent LLM intervention. In such a scenario, instead of giving all the previously gathered data to the LLM as a part of its intervention package, we supply it with the correlation matrix of the data instead. The correlation matrix summarizes the pairwise relationships between the input variables of the problem at hand, considerably reducing the input size for the LLM. In fact, for a dataset with $d$ input variables and $N>>d$ number of samples, the correlation matrix has only $d\times d$ elements instead of the raw dataset with $d\times N$ elements. Although the correlation matrix summarizes the data, the LLM loses the intricate details in the raw data. However, as a result of a condensed data summary, the LLM processes it much faster.